Circular No. 5829

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NO SUPERNOVAE 1993U AND 1993V A. V. Filippenko, T. Matheson, and L. C. Ho, University of California

at Berkeley, report that fully calibrated CCD spectra (range 320-1000 nm) obtained on June 28 UT with the Shane 3-m reflector at Lick Observatory reveal that SNe 1993U and 1993V (cf. IAUC 5818) are actually quasi-stellar objects (QSO). SN 1993U has a redshift of 0.591, based on the presence of broad (FWHM ∼ 4000 km/s) emission lines of Mg II (279.8 nm), Hγ, and Hβ superposed on a blue continuum. The emission-line redshift of the adjacent galaxy is 0.107, and there is a foreground M star next to it but diametrically opposite the QSO. SN 1993V has a probable redshift of 1.087, based on the presence of broad (FWHM ∼ 10 000 km/s) Mg II (279.8 nm) superposed on a red continuum. Weak C III] (190.9 nm) might also be visible. The redshift of the adjacent galaxy is 0.108, measured from narrow emission lines.

YY HERCULIS M. Friedjung, Institut d’Astrophysique, Paris, communicates: “As part

of a programme with C. C. Huang (Purple Mountain Observatory) involving observations of symbiotic stars between 810.0 and 870.0 nm, we have observed YY Her currently undergoing a bright phase (see IAUC 5819). The last observation made on 1993 June 30 can be compared with one made on 1992 June 15, all having been made with the 1.93-m telescope of the Haute Provence Observatory (+ CARELEC spectrograph; dispersion 3.3 nm/mm). The lines of the Ca II triplet (of which the two shorterwavelength ones at least were in absorption in 1992) now appear in emission, resolved from the nearby Paschen lines. The fluxes of the latter in emission on both dates, however, do not at first sight indicate very dramatic changes since 1992.”

NOVA AQUILAE 1993 Further photoelectric photometry (cf. IAUC 5818): June 23.197 UT,

V = 10.65 ± 0.02, B–V = +0.32 ± 0.04 (D. H. Kaiser, Columbus, IN); 24.93, V = 10.75 ± 0.02, B–V = +0.37 ± 0.03, U–B = –0.44 ± 0.05 (D. Hanzl, N. Copernicus Observatory, Brno; comparison star SAO 124317 = HD 178574); 26.05, V = 10.6 (H. Mikuž, Ljubljana, Slovenia); 28.90, V = 11.8 (Mikuž); 30.90, V = 12.4 (Mikuž).

1993 July 12 Daniel W. E. Green